# Efficiently Serving Large Multimodal Models Using EPD Disaggregation

### v2

@@ 4.1 SLO Attainment for End-to-End Generation
In this experiment, we evaluate the goodput of EPD and the baselines in an online setting, where 100 multimodal requests arrive following a Poisson process with rate $\lambda$. The length of output tokens is fixed to 10. We test requests with 2 and 4 images, each at a resolution of $4032\times 3024$. The details of the corresponding SLO criteria are discussed in Sec. B.3. The results for three models (columns) are presented in Fig. 5. The X-axis represents the per-GPU request rate ($\lambda$/#GPUs), and the Y-axis indicates the percentage of requests meeting both TTFT and TPOT requirements, with a 90% threshold denoted by a black dotted line.
@@
Figure 6: Distribution of TTFT (Y-axis) for various #Images/ Request (X-axis) for the MiniCPM-V 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right)
EPD outperforms all baselines, achieving over 90% SLO attainment at lower request rates. This is due to its ability to parallelize the computationally intensive image encoding step across multiple GPUs. DistServe and vLLM often maintain less than 10% attainment due to interference. Meanwhile, comparing the first and second rows for varying image counts per request (2 and 4), we observe that while more images per request increase workload, EPD maintains reasonable performance, whereas baseline SLO attainment drops significantly. Lastly, InternVL, being prefill-heavy, experiences queuing delays due to the higher number of image tokens, particularly at higher request rates. This effect exacerbates performance degradation from the 8B to the 26B model. In contrast, MiniCPM-V, optimized to generate fewer image tokens, avoids these delays and achieves better overall latency. Further, we provide more results with 6 and 8 images in Sec. D.1.
Next, we repeat the experiment using the non-synthetic video question-answering dataset, NextQA [18]. For this, we randomly sampled 100 examples, with input text token lengths ranging from 4 to 21 (average: 11.42) and output token lengths ranging from 1 to 7 (average: 2.75). Each video request was represented by 8 uniformly sampled frames, and we used MiniCPM-V 2.6 in the experiment, adhering to the SLO criteria of TTFT = 5.60 and TPOT = 0.06. As illustrated in Fig. 7, EPD is the only framework achieving 90% SLO attainment at low request rates, demonstrating its superior effectiveness in handling diverse, real-world-like workloads compared to DistServe and vLLM.
Figure 7: SLO attainment for end-to-end inference on the NextQA dataset. As seen, EPD outperforms other baselines by a significant margin.
@@ 4.2 First Token Generation Latency
Multimodal requests often place a heavy load on both the E and P phases, making the first token generation a key latency bottleneck. Therefore, in this experiment, we analyze the first token latency for various baselines. The results presented in Fig. 6 shows box plots of TTFT distributions for three different models. Note that, as the decoding phase is excluded, the vLLM baseline is equivalent to DistServe and is thus skipped. Requests are generated according to a Poisson distribution with a fixed request rate $\lambda$. Specifically, $\lambda=0.25$ for the MiniCPM-V 2.6, and $\lambda=0.08$ for both the InternVL2-8B and InternVL2-26B. Thanks to intra-request parallelization, EPD significantly outperforms both baselines, EPD and DistServe. Specifically, TTFT is reduced by up to 71.9%, 32.8%, and 44.9% compared to the DistServe baseline for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models, respectively. The performance gain grows with the number of images per request, demonstrating EPD’s efficiency in handling heavy multimodal workloads.
@@ 4.3 Memory Savings through Stage Disaggregation
In this section we analyze the memory savings achieved by disaggregating the encoding and prefill stages. The E workers can save memory as they do not require LLM weights or the KV cache. Analyzing only the weight size indicates memory reduction of approximately 95%, 96.2%, and 78.3% for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models. Similarly, for the P workers, memory saving of about 5%, 3.7%, and 21.6% are achieved, respectively. In practice, since KV cache is also not required at E workers, the memory saving can be even higher (93.3% saving, i.e., 15$\times$ lower, according to profiling). These reductions in memory usage enable the EPD system to support higher numbers of images per request and batch sizes (discussed in the following experiments) and larger KV cache sizes (discussed in Appendix D.2).
@@
Table 1: Comparison of maximum supported number of images/ request. Batch size is fixed to 1. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
EPD Supports a Higher Number of Images per Request: We compare the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B. The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and with 80% of the available memory allocated for KV cache. As shown in Tab. 1, EPD handles more images per request than DistServe, for example, at 4032$\times$3024 resolution, 7$\times$ more for InternVL2-26B and 10$\times$ more for InternVL2-8B. For InternVL2-8B, the limit of 19 images per request is due to its maximum context length. Without this constraint, a larger number of images per request could be supported.
EPD Supports Higher Batch Sizes: We compare maximum supported batch sizes for E and P stages across different settings of image resolution and models in Tab. 2. As seen, EPD significantly outperforms DistServe for both E and P batch sizes. For example, with InternVL2-26B at 787$\times$444 resolution, EPD supports a batch size of 22 for encoding vs. DistServe’s 1, achieving 22$\times$ improvement. Similarly, for MiniCPM-V 2.6 at 787$\times$444, EPD achieves a batch size of 29 vs. DistServe’s 2 for prefill, a 14.5$\times$ improvement.
@@
Table 2: Comparison of maximum supported E and P batch size. #images/ request is fixed to 10. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
@@ 4.4 Ablation Study
Effect of IRP: We analyze the effect of ablating IRP, presented in Section 3.2.2. The experimental settings are same as TTFT experiment in Sec. 4.2. As shown in Tab. 3, removing the IRP feature negatively affects TTFT across various images/ request. Moreover, the degradation exacerbates as the number of images/ request increases. This is because the IRP feature allows parallelization of encoding load within the same request across multiple GPUs.
@@
Table 4: Effect of Offline Optimizer. When the configuration is set randomly, the goodput is 2.2× worse on average. $\downarrow$ indicates that lower is better and $\uparrow$ means that larger is better.
Effect of Offline Optimizer: By default, the EPD system collects workload samples and finds the optimal config offline using the optimizer. To demonstrate its effect, we conduct an experiment without using the optimizer and set configs randomly. Specifically, we uniformly sample 10 configs at random and report the estimated expectation of the performance metric in the second row of Tab. 4. For a fair comparison, when evaluating the TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which is the goodput of the EPD system. (See Appendix B.4 for more details) As shown in Tab. 4, disabling the optimizer causes significant performance degradation, in terms of both goodput and TTFT. This demonstrates the necessity of the optimizer, especially when the EPD system has many tunable configs.
Effect of Dynamic Role Switching: To analyze the impact of dynamic role switching, we conduct a controlled experiment simulating a shift in workload characteristics. Specifically, we generate 100 requests following the same static workload configuration as described in Section 3.2.4. However, to introduce an artificial workload change, we set the first 10 requests to generate 50 output tokens, while the remaining 90 requests generate 500 output tokens. The request arrival rate is fixed at 3 requests per second. The results are shown in Tab. 5.
@@
Table 5: Effect of ablating dynamic role switching feature from EPD. Disabling the worker migration feature negatively affects the TPOT (2.4x worse), which results in significantly poor end-to-end latency (2.2x worse). Results are averaged over 100 requests with one 4K image per request. $\downarrow$ indicates that lower is better.
Without dynamic worker migration, the system performs poorly because it remains locked in the initial configuration (5E1P2D, optimized offline for 50 tokens) and is unable to adapt to the increased decoding demand. In contrast, the EPD system with migration dynamically reconfigures itself (2E1P5D) to handle the new workload (500 output tokens) by shifting 3 E instances to D, resulting in approximately 2× better performance.
@@ 4.5 Adaptation to Neural Processing Units (NPUs)
This section presents the results of adapting the proposed EPD framework to Huawei Ascend NPUs. For more details on the implementation and experiments, see Sec. B.5.
First, we compare the EPD-NPU against the vLLM and DistServe baselines in terms of SLO attainment using the InternVL2-8B LMM. We used the same settings as mentioned in Sec. 4.1, except that we used a heavy encoding workload of 8 4032x3024 images per request. The optimal configuration for this workload was found to be 5E2P1D and the corresponding results are presented in Figure 8. As seen, EPD is the only setting that can achieve the SLO requirements, while the other baselines are completely decimated even for low request rates.
Figure 8: SLO curve for InternVL2-8B with the following requirements (TTFT=8.5 and TPOT=0.12) on the controlled experimental setup with 8 images per request.
Next, we compare the TTFT of EPD-NPU against vLLM. We found that EPD-NPU demonstrated a 35.2% improvement over vLLM, which is significantly more than the 24.4% improvement for the GPU scenario. We attribute this additional $\sim$10to the proportionally heavier encode of LLMs on NPU compared to GPU. Please refer to Sec. D.3 for more details. This results in greater benefits of disaggregating the encode from prefill on NPU in comparison to GPU.
@@ 5 Conclusion
In this paper, we present a novel approach to optimizing Large Multimodal Model (LMM) systems through the disaggregation of key processing stages. By separating the encoding, prefill, and decoding tasks into distinct stages, our system offers enhanced flexibility in resource allocation, enabling more efficient management of computational and memory resources. This disaggregation, combined with dynamic resource allocation, asynchronous token transfer, and advanced parallelization strategies, directly addresses several critical challenges in LMM deployment, including latency reduction, memory optimization, and efficient computational resource usage.
Appendix A Optimizer Details
We recapitulate the configuration optimization problem:
  • Scheduling involves two main decisions: First, which workers should each request be assigned to? Second, how should the order of requests be determined within a worker queue? To solve these decisions, we adapt strategies from the DistServe framework [21]. In the encoding stage, when a request arrives, it is assigned and pushed to an instance queue. Between different stages, global queues are used, and each available engine pulls proactively from the queues. Possible assignment strategies in the encoding stage include Round-Robin or Least-Loaded First for assigning requests. Once a request is assigned to a worker’s queue, we can apply ordering strategies like first-come-first-serve or shortest-job-first, or more complex strategies that prioritize requests based on their Service Level Objectives (SLOs). For simplicity, we constrain that all instances within the same stage share the same scheduling strategy.
Appendix B Implementation Details
B.1 Hyper-parameters in EPD system
B.2 LMM Details
MiniCPM-V 2.6 integrates a SigLip-400M vision encoder, comprising 400 million parameters, with a Qwen2-7B language model, containing 7.6 billion parameters, culminating in 8 billion parameters. It excels in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains [19].
B.3 SLO Criteria
Tab. 6 outlines the SLO criteria for TTFT and TPOT across various models and different numbers of images per request (#I/R). These criteria are empirically derived based on the characteristics of the underlying models, such as the computational complexity of the MME and LLM. We also consider what is realistically achievable by both our method and the baselines on a fixed number (8 GPUs) used in experiments. Across models, as #I/R increases, there is an approximately linear increase in the TTFT criteria due to the higher encoding load and additional tokens generated during the prefill stage. In contrast, TPOT requires only minor adjustments since it is not directly impacted by changes in #I/R.
MiniCPMv 2.6 8B InternVL 8B InternVL 26B
#I/R TTFT TPOT TTFT TPOT TTFT TPOT
2 1.40 0.04 1.20 0.05 3.50 0.07
4 2.60 0.04 2.40 0.06 7.05 0.08
6 3.90 0.06 3.55 0.09 11.00 0.95
8 5.10 0.06 5.00 0.18 15.00 0.15
Table 6: Requirements for TTFT and TPOT (in seconds) used for SLO computation across various models and number of images per request (#I/R) in the experiments.
B.4 Details for the Ablation Experiment of Offline Optimizer
B.5 NPU Implementation Details
There are intrinsic differences between NPU and GPU architectures and processing characteristics. GPUs are general-purpose compute resources whereas NPUs have been specifically designed for neural networks. NPUs are optimized for FP16 calculations whereas GPUs are optimized for FP32 calculations. As stated above, CANN operators are used on Ascend NPU hardware in place of CUDA operators. These differences between CUDA and corresponding CANN operators, having been optimized for the hardware they are built upon, result in different latencies and trade-offs in inference as well as communication between NPU/GPU chips. This results in different encode/prefill/decode latencies which may be present different benefits from EPD and IRP.
The EPD-NPU system was implemented in a similar manner to our EPD-GPU system as described in Appendix B. All EPD-NPU experiments utilize IRP. The EPD-NPU distributed serving system was developed and deployed on a cluster of 910B3 NPUs, each with 64GB of high bandwidth memory. Some primary differences between the two implementations include:
Figure 9: Left: The impact of tuning the number of encoding workers in the EPD method. The annotation ‘xEyP’ indicates that x encoders and y prefill workers are allocated for EPD, and ‘7P’ means that the DistServe always allocates 7 P workers for the encoding and prefill stage requests. Middle: The impact of the number of images per request in the workload. Right: The impact of encoding and prefill batch sizes in the EPD method (encoding and prefill batch sizes are set to be equal).
These differences in characteristics provide an opportunity to investigate the generalizability of our EPD framework and further understand the optimal setting/environment for disaggregating encode and prefill.
For EPD-NPU, we use eight 910B3 NPUs and evaluate the system by running the controlled experimental setup (as described in the experiments section of the main paper) in an online setting. Specifically, the number of input tokens was 22, length of output tokens was 10 with early stopping was disabled. The SLO requirements were also selected in a similar manner as the GPU experiments. The requirements for InternVL2-8B with 8 images per request were 8.5 and 0.12 for TTFT and TPOT, respectively. Each server had 192 CPUs and 1.5TB of RAM. The CANN version was 7.6. The maximum model length was set to 27000, KV Cache utilization was set to 83% And the maximum number of multimedia data of 8 was imposed. EPD-NPU also used a version of Ascend-vLLM (version 0.6.3.post1) built upon a newer version of vLLM than the one used in EPD-GPU. Unless otherwise stated, the hyper-parameters are identical to those described in B.1.
Appendix C Throughput in Offline Settings
In this section, we demonstrate the benefits of the EPD method in terms of throughput, specifically focusing on its memory efficiency and the ability to support higher batch sizes for both the E and P stages.
Consider an offline scenario where a batch of requests is submitted, allowing the system to process them overnight with a focus on maximizing end-to-end (E2E) throughput. This is an additional performance metric complementing the main body of the paper, measured in terms of requests completed per second and the system’s ability to handle high request volumes in offline settings. In the traditional DistServe method, memory is allocated for both the encoding and prefill stages on the same worker, which is constrained by the total memory capacity of the GPU.
This memory limitation becomes especially significant in heterogeneous environments with low-end GPUs that have limited memory. In these cases, the DistServe method struggles to fully utilize low-end GPUs, as the combined memory demands of the encoding and prefill stages exceed the available memory. In contrast, our disaggregated EPD method is more cost-effective, as it assigns the low-end GPUs exclusively to the encoding stage, allowing better utilization of available memory. In this experiment, we consider a less extreme scenario where the memory demands of the prefill workers in the DistServe method do not exceed the memory capacity of the low-end GPUs, allowing for a batch size of 1 for both the encoding and prefill stages.
(a) MiniCPMv-2.6                            (b) InternVL2-8B                        (c) InternVL2-26B
Figure 10: Columns (a), (b), and (c) correspond to MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively. The top and bottom rows show results for 6 and 8 images per request, respectively. As expected, EPD maintains a significant performance advantage over all baselines, demonstrating robustness of the system as the image count increases.
We use 8 A800 GPUs for this experiment. For EPD, we allocate 5 GPUs for encoding (E), 2 GPUs for prefill (P), and 1 GPU for decoding (D), with the maximum batch sizes set to 8, 8, and 128 for E, P, and D, respectively. In the DistServe method, we allocate 7 GPUs for prefill (P) and 1 GPU for decoding (D), with the maximum batch sizes of 1 and 128, respectively. The workload consists of 1,000 requests, each with a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens. We tune the system’s hyperparameters as outlined in Sec. 3.2.3 and evaluate their impact on throughput. The results are presented in Fig. 9.
The left plot illustrates the importance of selecting an appropriate GPU configuration. As described in Sec. 3.2.3, our algorithm automatically identifies that a 5E2P configuration (5 encoder and 2 prefill workers) maximizes E2E throughput. The middle plot shows that EPD achieves superior throughput when the number of images per request is low, demonstrating that EPD’s architecture prevents the encoder from becoming a compute-bound bottleneck. The right plot highlights that EPD’s performance is relatively insensitive to encoding and decoding batch sizes, consistently outperforming the DistServe method. This flexibility allows us to either choose a relatively large batch size or rely on the algorithm in Sec. 3.2.3 to automatically select an optimal batch size.
Appendix D Additional Experiments
D.1 SLO Attainment for End-to-End Generation for 6 and 8 images
In this section, we show additional results from the experiment in Sec. 4.1 pertaining to 6 and 8 images per request. As seen, the SLO attainment for EPD begins to decline as number of images rise, particularly at higher request rates. This decline reflects the increasing computational demand associated with processing multiple high-resolution images within a single request. Nevertheless, EPD continues to outperform all baselines, which struggle even at low request rates and fail to scale effectively under increased workloads.

### v3

@@ 4.1 SLO Attainment for End-to-End Generation
In this experiment, we evaluate the goodput of EPD and the baselines in an online setting where 100 multimodal requests arrive following a Poisson process with rate $\lambda$. The length of output tokens is fixed to 10. We test requests with 2 and 4 images, each at a resolution of $4032\times 3024$. The details of the corresponding SLO criteria are discussed in Section E.3. The results for three models (columns) are presented in Figure 5. The X-axis represents the overall request rate ($\lambda$), and the Y-axis indicates the percentage of requests meeting both TTFT and TPOT requirements, with a 90% threshold denoted by a black dotted line.
@@
Figure 6: Distribution of TTFT (Y-axis) across varying numbers of images per request (X-axis) for (a) MiniCPM-V 2.6, (b) InternVL2-8B, and (c) InternVL2-26B. Each plot illustrates the latency behavior under increasing input sizes.
EPD outperforms all baselines, achieving over 90% SLO attainment at lower request rates. This is due to its ability to parallelize the computationally intensive image encoding step across multiple GPUs. DistServe and vLLM often maintain less than 10% SLO attainment due to interference. Meanwhile, comparing the first and second rows for varying image counts per request (2 and 4), we observe that, while more images per request increase the workload, EPD maintains reasonable performance, whereas baselines’ SLO attainment drops significantly. Lastly, InternVL, which is prefill-heavy, experiences queuing delays due to the higher number of image tokens, particularly at higher request rates. This effect further exacerbates the performance degradation from the 8B to the 26B model. In contrast, MiniCPM-V, optimized to generate fewer image tokens, avoids these delays and achieves lower overall latency. Further results with 6 and 8 images are provided in Section A.4.
Next, we repeat the experiment using the non-synthetic video question-answering dataset, NextQA (Xiao et al., 2021). To do so, we randomly sampled 100 examples, with input text token lengths ranging from 4 to 21 (average: 11.42) and output token lengths ranging from 1 to 7 (average: 2.75). Each video request was represented by 8 uniformly sampled frames, and we used MiniCPM-V 2.6 in the experiment, adhering to the SLO criteria of TTFT = 5.60 and TPOT = 0.06. As illustrated in Figure 7, EPD is the only framework achieving 90% SLO attainment at low request rates, demonstrating its superior ability to handle real-world workloads compared to DistServe and vLLM.
Figure 7: SLO attainment ($\uparrow$) versus request rate on the NextQA dataset using the MiniCPM-V 2.6 model. EPD consistently achieves higher SLO attainment compared to all baselines.
Finally, we conduct the same experiment using the Video-MME (Fu et al., 2024) dataset. We evaluate SLO attainment, defined as TTFT $\leq$ 3.1s and TPOT $\leq$ 0.025s, on 100 randomly sampled Video-MME examples using MiniCPM-V 2.6. Each video is represented by 64 uniformly sampled frames, following the MiniCPM frame configuration reported on the Video-MME leaderboard. The results are shown in Figure 8.
As seen, EPD consistently outperforms vLLM and DistServe across all rates, demonstrating a strong generalization to temporal multimodal workloads, such as videos.
Figure 8: SLO attainment ($\uparrow$) versus request rate on the Video-MME dataset using the MiniCPM-V 2.6 model. EPD significantly outperforms competing baselines across all request rates.
Method #Frames 8 16 32 64
vLLM 0.42 0.82 1.59 3.11
DistServe 0.42 0.81 1.54 3.08
EPD (ours) 0.24 0.30 0.49 1.00
Table 1: Mean TTFT latency (in seconds) ($\downarrow$) for varying video lengths at a fixed request rate of 1 request/sec. Results are averaged over 100 Video-MME samples. EPD achieves the lowest latency across all video lengths.
@@ 4.2 First Token Generation Latency
Multimodal requests often impose a heavy load on both the E and P phases, making the first token generation a key latency bottleneck. Therefore, in this experiment, we analyze the first token latency for various baselines. The results presented in Figure 6 show box plots of TTFT distributions for three different models. Note that, as the decoding phase is excluded, the vLLM baseline is equivalent to DistServe and is thus omitted. Requests are generated according to a Poisson distribution with a fixed request rate $\lambda$. Specifically, $\lambda=0.25$ for MiniCPM-V 2.6 and $\lambda=0.08$ for both InternVL2-8B and InternVL2-26B. Thanks to intra-request parallelization, EPD significantly outperforms both vLLM and DistServe. Specifically, TTFT is reduced by up to 71.9%, 32.8%, and 44.9% compared to the DistServe baseline for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
Additionally, we conducted a TTFT comparison on the Video-MME (Fu et al., 2024) dataset for various baselines. The results are presented in Table 1. As seen, EPD achieves a latency reduction of up to 68.2% over DistServe and up to 69.2% over vLLM. Moreover, the performance gap widens with increasing video length—for instance, at 8 frames, EPD reduces latency by 42.9%, and at 64 frames, the reduction reaches 67.5% relative to DistServe. These results highlight EPD’s superior scalability and robustness under increasingly demanding video processing workloads.
@@ 4.3 Memory Savings through Stage Disaggregation
In this section, we analyze the memory savings achieved by disaggregating the encoding and prefill stages. The E workers can save memory as they do not require LLM weights or the KV cache. Analyzing only the weight size indicates memory reduction of approximately 95%, 96.2%, and 78.3% for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models. Similarly, for the P workers, memory savings of about 5%, 3.7%, and 21.6% are achieved, respectively. In practice, since KV cache is also not required at E workers, the memory saving can be even higher (93.3% saving, i.e., 15$\times$ lower, according to profiling). These reductions in memory usage enable the EPD system to support higher numbers of images per request and batch sizes (discussed in the following experiments) and larger KV cache sizes (discussed in Appendix A.2).
@@
Table 2: Comparison of the maximum number of images supported per request for various image resolutions across different models. Higher values are better; best values in each row are italicized.
EPD Supports a Higher Number of Images per Request: We compare the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B. The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and with 80% of the available memory allocated for the KV cache. As shown in Table 2, EPD handles more images per request than DistServe, for example, at 4032$\times$3024 resolution, 7$\times$ more for InternVL2-26B and 10$\times$ more for InternVL2-8B. For InternVL2-8B, the limit of 19 images per request is due to its maximum context length. Without this constraint, a larger number of images per request could be supported.
EPD Supports Higher Batch Sizes: We compare maximum supported batch sizes for E and P stages across different settings of image resolution and models in Table 3. The number of images per request is fixed to 10, and the KV cache is allocated to utilize 80% of the available free memory. As seen, EPD significantly outperforms DistServe for both E and P batch sizes. For example, with InternVL2-26B at 787$\times$444 resolution, EPD supports a batch size of 22 for encoding vs. DistServe’s 1, achieving 22$\times$ improvement. Similarly, for MiniCPM-V 2.6 at 787$\times$444, EPD achieves a batch size of 29 vs. DistServe’s 2 for prefill, a 14.5$\times$ improvement.
@@
Table 3: Comparison of the maximum supported batch sizes for E and P stages across different models and image resolutions. Higher values are better; italicized values indicate the best in each row. OOM denotes cases where the model ran out of memory.
@@ 4.4 Ablation Study
Effect of IRP: We analyze the effect of ablating IRP, presented in Section 3.2.2. The experimental settings are the same as the TTFT experiment in Section 4.2. As shown in Table 4, removing the IRP feature negatively affects TTFT across various images/ request. Moreover, the degradation exacerbates as the number of images/ request increases. This is because the IRP feature allows parallelization of encoding load within the same request across multiple GPUs.
@@
Table 5: Ablating the offline optimizer reduces goodput by 2.2× on average when configurations are selected randomly. $\downarrow$ indicates lower is better; $\uparrow$ indicates higher is better.
Effect of Offline Optimizer: By default, the EPD system collects workload samples and finds the optimal configuration offline using the optimizer. To demonstrate its effect, we conduct an experiment without using the optimizer and select configurations randomly. Specifically, we uniformly sample 10 configurations at random and report the expected value of the performance metric in the second row of Table 5. For a fair comparison, when evaluating TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which corresponds to the goodput of the EPD system (see Appendix E.4 for more details). As shown in Table 5, disabling the optimizer causes significant performance degradation in terms of both goodput and TTFT. This highlights the importance of the optimizer, especially when the EPD system has many tunable configurations.
Effect of Dynamic Role Switching: To analyze the impact of dynamic role switching, we conduct a controlled experiment simulating a shift in workload characteristics. Specifically, we generate 100 requests following the same static workload configuration as described in Section 3.2.4. However, to introduce an artificial workload change, the first 10 requests generate 50 output tokens, while the remaining 90 requests generate 500 output tokens. The request arrival rate is fixed at 3 requests per second. The results are shown in Table 6.
@@
Table 6: Ablating dynamic role-switching from EPD degrades TPOT by 2.4× and increases end-to-end latency by 2.2×. Results are averaged over 100 requests with one 4K image each. $\downarrow$ indicates lower is better.
Without dynamic worker migration, the system performs poorly because it remains fixed in the initial configuration (5E1P2D, optimized offline for 50 tokens) and is unable to adapt to the increased decoding demand. In contrast, the EPD system with migration dynamically reconfigures itself (2E1P5D) to handle the new workload (500 output tokens) by shifting three E instances to D, resulting in approximately 2× better performance.
@@ 4.5 Adaptation to Neural Processing Units (NPUs)
We present the results of adapting the proposed EPD framework to Huawei Ascend NPUs. For more details on the implementation and experiments, see Appendix F.
First, we compare EPD-NPU against the vLLM and DistServe baselines on the InternVL2-8B LMM. We evaluate SLO attainment, defined as TTFT $\leq$ 8.5s and TPOT $\leq$ 0.12s on eight 4K images per request. We used the same settings as described in Section 4.1, except that we employed a heavy encoding workload of eight $4032\times 3024$ images per request. The optimal configuration for this workload was found to be 5E2P1D, and the corresponding results are presented in Figure 9. As shown, EPD is the only configuration that achieves the SLO requirements, while the other baselines fail to meet the SLOs entirely, even at low request rates.
Figure 9: SLO attainment ($\uparrow$) on NPUs under varying request rates for a synthetic workload using the InternVL2-8B model. EPD maintains positive SLO attainment under strict TTFT constraints, while baselines fail to meet the criteria.
Next, we compare the TTFT of EPD-NPU against vLLM. EPD-NPU achieves a 35.2% improvement over vLLM, surpassing the 24.4% improvement observed in the GPU scenario. This additional $\sim$10improvement is attributed to the proportionally heavier encoding workload of LLMs on NPUs compared to GPUs. Please refer to Section F.1 for more details. As a result, disaggregating the encoding from the prefill phase yields greater benefits on NPUs than on GPUs.
@@ 5 Conclusion
In this paper, we proposed a novel approach for optimizing LMM systems via disaggregation of key processing stages. By separating the encoding, prefill, and decoding phases into distinct stages, our system provides greater flexibility in resource allocation, enabling more efficient management of computational and memory resources. This disaggregation, together with dynamic resource allocation, asynchronous token transfer, and advanced parallelization strategies, directly addresses several critical challenges in LMM deployment, including latency reduction, memory optimization, and efficient computational resource usage. We validated the effectiveness of our system through extensive experiments. Finally, we outline the system’s limitations and future directions in Appendix B and Appendix C, respectively.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Experiments
A.1 Adaptation to Audio modality
In this section, we extend EPD to the audio modality to evaluate its effectiveness beyond vision tasks, highlighting the method’s generalizability. Specifically, we evaluate EPD in an online, audio setting using the ultravox-v0_3 model, based on LLaMA3.1-8B. To simulate an encode-intensive workload, each request is configured to include 24 audio files. Evaluation metrics include SLO attainment—defined as TTFT $\leq$ 2.0 s and TPOT $\leq$ 0.025 s—and goodput (requests per second, r/s). As shown in Table 7, EPD consistently outperforms vLLM and DistServe across all request rates, demonstrating superior efficiency and reliability. Combined with prior results in the video domain, these findings reinforce EPD’s robustness and generalizability across modalities.
Rate (r/s) SLO Attainment Rates ($\uparrow$)
vLLM DistServe EPD
0.10 0.99 0.99 0.99
0.25 1.00 0.94 0.99
0.50 0.99 0.89 1.00
1.00 0.91 0.72 0.96
1.10 0.87 0.69 0.93
1.15 0.87 0.68 0.93
Goodput (r/s) $\uparrow$ 1.01 0.45 1.16
Table 7: SLO attainment results ($\uparrow$) for online audio benchmarking with ultravox-v0_3 (24 audio files per request). All baselines use 4 GPUs: vLLM operates in data-parallel (DP) mode, DistServe uses a 3P1D configuration, and EPD adopts a 2E1P1D setup. EPD achieves consistently high SLO attainment and the highest goodput across all request rates.
A.2 EPD Supports Larger KV Cache Sizes
In this section, we compare the maximum KV cache size that can be allocated across various baselines. The batch size is fixed at 1, and the image resolution is set to 4032$\times$3024. As shown in Table 8, EPD supports significantly larger KV cache sizes than DistServe. For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for DistServe, representing a 2.2$\times$ improvement. Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an OOM error, indicating that it cannot process 20 images even with a KV cache size of 0. Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.